OpenStudy (erikaxx):

what is the integral of e^2y/x dx

OpenStudy (erikaxx):

\[e ^{2(\frac{ y }{ x })} dx\]

OpenStudy (erikaxx):

is it e^2lnx

ganeshie8 (ganeshie8):

Looks wrong. You cannot integrate y with respect to x w/o plugging in the expression for y

ganeshie8 (ganeshie8):

May I know how you arrived at that integral ?

OpenStudy (erikaxx):

ah im applying it in linear equation

ganeshie8 (ganeshie8):

May I see your work ? if possible take a screenshot of your work and attach here

OpenStudy (erikaxx):

okay

OpenStudy (erikaxx):

here @ganeshie8

ganeshie8 (ganeshie8):

nice there is a mistake in third line

ganeshie8 (ganeshie8):

you have flipped the sign of 2y/x

OpenStudy (erikaxx):

yes becauseof the formula

ganeshie8 (ganeshie8):

What formula

ganeshie8 (ganeshie8):

There are no formulas in math that allow you to flip the signs magically

OpenStudy (erikaxx):

here in linear

ganeshie8 (ganeshie8):

Given eqn : ``` xdy/dx - 2y = x^3cos4x ```

ganeshie8 (ganeshie8):

dividing x through out gives ``` dy/dx - 2y/x = x^2cos4x ```

ganeshie8 (ganeshie8):

Notice that `P(x) = -2/x`

OpenStudy (erikaxx):

yes, but the px is -2 y/x

ganeshie8 (ganeshie8):

No P(x) is simply -2/x

OpenStudy (erikaxx):

huh, how come. the equation says it is `2y/x

ganeshie8 (ganeshie8):

I want you take some time and compare above equation with the general linear form

OpenStudy (erikaxx):

ah got it the y will be cancel

ganeshie8 (ganeshie8):

Compare below two eqns : ``` dy/dx - 2y/x = x^2cos4x ``` and ``` dy/dx + P(x)y = Q(x) ```

OpenStudy (erikaxx):

yes got it

ganeshie8 (ganeshie8):

P(x) = -2/x you don't need to wry about Q(x)

OpenStudy (erikaxx):

ah yes, the proceed to I.F

ganeshie8 (ganeshie8):

so the integrating factor is just \[IF = e^{\int -2/x\,dx}\]

OpenStudy (erikaxx):

YES so e^2lnx

ganeshie8 (ganeshie8):

careful with the sings

OpenStudy (erikaxx):

with negative

ganeshie8 (ganeshie8):

its e^(-2lnx)

OpenStudy (erikaxx):

yes so it is now -2x

ganeshie8 (ganeshie8):

How ?

OpenStudy (erikaxx):

e^ln is 1 so then -2x

ganeshie8 (ganeshie8):

No

OpenStudy (erikaxx):

huh

ganeshie8 (ganeshie8):

\[e^{-2\ln x} = \left(e^{\ln x}\right)^{-2} = (x)^{-2} = \dfrac{1}{x^2}\]

OpenStudy (erikaxx):

ah, i know now. then in solution

ganeshie8 (ganeshie8):

recall that \((x^m)^n = x^{mn}\)

OpenStudy (erikaxx):

(I,F)y = int Q(x) (I.F) dx +c

OpenStudy (erikaxx):

\[(\frac{ 1 }{ x ^{2} } )y = \int\limits x ^{2} \cos 4x (x)(\frac{ 1 }{ x ^{2} })\]

ganeshie8 (ganeshie8):

\[(\frac{ 1 }{ x ^{2} } )y = \int\limits x ^{2} \cos (4x)(\frac{ 1 }{ x ^{2} })\]

OpenStudy (erikaxx):

with plus c

ganeshie8 (ganeshie8):

\[(\frac{ 1 }{ x ^{2} } )y = \int\limits x ^{2} \cos (4x)(\frac{ 1 }{ x ^{2} })\,dx\]

ganeshie8 (ganeshie8):

you can put +c after evaluating the integral

OpenStudy (erikaxx):

ah okay \[(\frac{ 1 }{ x ^{2} }) y = \int\limits \frac{ x ^{3} }{ 3 } \sin 4x \]

ganeshie8 (ganeshie8):

Don't put the integral sign after evaluating the integral

ganeshie8 (ganeshie8):

also that is wrong

OpenStudy (erikaxx):

oops sorry

ganeshie8 (ganeshie8):

\[(\frac{ 1 }{ x ^{2} } )y = \int\limits x ^{2} \cos (4x)(\frac{ 1 }{ x ^{2} })\,dx\] right hand, cancel x^2 top and bottom first

OpenStudy (erikaxx):

ah okay so \[\frac{ 1 }{ x^2 } y = \sin4x + c\]

OpenStudy (erikaxx):

then multiply both side by x^2

ganeshie8 (ganeshie8):

Still wrong

ganeshie8 (ganeshie8):

whats the integral of cos(4x) ?

OpenStudy (erikaxx):

1/4 sin 4x

ganeshie8 (ganeshie8):

Yes, 1/4 is missing in your answer

OpenStudy (erikaxx):

yep i forgot tho rewrite \[\frac{ 1 }{ x^2 } y = \frac{ 1 }{ 4 } \sin 4x + c\]

OpenStudy (erikaxx):

then multiply both side by 4x^2

OpenStudy (erikaxx):

so it become \[4y = x^2 \sin 4x + 4cx^2\]

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (erikaxx):

thanks for the help. need to know more about linear.

OpenStudy (erikaxx):

okay thanks i looked for it

ganeshie8 (ganeshie8):

np :)